OpenStudy (anonymous):

the endpoints of circle q's diameter are (2,3) and (-1,8). Find the standard form of the equation of Circle Q

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

OK , so now you can find the length of the diameter and then find the distance of the radius , do u know the distance formula?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

use it and tell me ur answer for the diameter

OpenStudy (anonymous):

the square root of 34.. which equals to about 5.8

OpenStudy (anonymous):

leave it as root34 for now , so radius is \[\sqrt{34}/2\]

OpenStudy (anonymous):

now use the formula \[r ^{2} = \left( x - a \right)^{2} + \left( y - b \right)^{2}\]

OpenStudy (anonymous):

r = root34 / 2 a = 2 b = 3

OpenStudy (anonymous):

how did you get 2 and 3?

OpenStudy (anonymous):

oh ,umm ya hold up , first you need to find the midpoint of the line ( the center) formula : (x , y) = (x1 + x2)/2 , (y1 + y2)/2 yse the points u have for x1,x2,y1,y2

OpenStudy (anonymous):

*use

OpenStudy (anonymous):

so the midpt would be (1/2,11/2)?

OpenStudy (anonymous):

exactly , now use those for a & b values in the other equation

OpenStudy (anonymous):

Thank you so much!!!